Consider the reaction MnO2 + 4HCl → MnCl2 + Cl2 + 2H2O. If the both the


reactants have a mass of 100 grams, which of the two will be used up first? How


many grams of MnCl2 will be produced after the reaction? (MnO2 =


86.94 g/mol, HCl = 36.46 g/mol, MnCl2 = 125.8 g/mol)


1
Expert's answer
2022-04-29T10:57:02-0400

From the equation 1 mole of MnO2 reacts with 4 moles of HCl

That means that "86.94 g" of MnO2 react with "4*36.46=145.84 g" of HCl.

Since we have 100 g of MnO2, it needs more than 100 g of HCl.

So HCl will be used up first.

Find the mass of MnCl2:

"m(MnCl_2) = m(HCl)*M(MnCl_2)*\nu(MnCl_2)/M(HCl)*\nu(HCl) = 100*125.8*1/36.46*4 = 86.26 (g)"

Answer: HCl will be used up first. 86.26 g of MnCl2 will be produced after the reaction.